### v4

# Canonical bundle formula and degenerating families of volume forms

Abstract.
General canonical bundle formula due to Kawamata and others has played fundamental roles in algebraic geometry. We show that the canonical bundle formula has analytic characterization in terms of fiberwise integration, which confirms a folklore conjecture. The proof uses $L^{2}$ metrics and the valuative equivalence of plurisubharmonic singularities. As an application, we identify the singularity of the Ohsawa measure from a general $L^{2}$ extension theorem of Demailly for log canonical pairs, in terms of the singularity in Kawamata’s subadjunction. As another application, we give a partial answer to a question of Berndtsson on semipositivity theorems.
Key words and phrases:
Canonical bundle formula, $L^{2}$ metrics, $L^{2}$ extension theorems, semipositivity theorems, fiberwise integration, plurisubharmonic functions
2010 Mathematics Subject Classification:
14D06, 32J25, 32U05, 14E30
1. Introduction
General canonical bundle formula [Ka98] (cf. [FM00], [Am04], [Ko07]) has played fundamental roles in higher dimensional algebraic geometry. The prototype of such formula arises from the classical work of Kodaira on elliptic fibrations [K64, Thm. 12]. The setting of the general canonical bundle formula and our main results is as follows.
1.1. Statements of the main results
Let $f:X\to Y$ be a surjective projective morphism with connected fibers between complex manifolds. We will say that $f:(X,R)\to(Y,B)$ satisfies Kawamata’s condition [Ka98] (see Definition 4.1) when $R$ is a simple normal crossing (or snc) divisor on $X$ and $B$ is a reduced snc divisor on $Y$ such that $f$ is smooth over $Y\setminus B$, $\operatorname{red}(R)+f^{*}B$ is an snc divisor on $X$ and the vertical parts of $R$ maps into $B$. Moreover, the horizontal parts of $R$ satisfy the conditions in Definition 4.1.
Let $L:=\mathcal{O}(R)$ be the $\mathbb{Q}$-line bundle associated to $R$. Let $M$ be the $\mathbb{Q}$-line bundle on $Y$ defined by the relation $K_{X}+L=f^{*}(K_{Y}+M)$.  [^1] See conventions in the beginning of Section 2 for our notation of hermitian line bundles. When $M$ is a $\mathbb{Z}$-line bundle (i.e. a usual holomorphic line bundle), it is equal to the direct image $M=f_{*}(K_{X/Y}+L)$ for which the $L^{2}$ metric is defined in Proposition 3.6. Our first main result is the following metric version of the canonical bundle formula [Ka98].
Theorem 1.1.
Let $f:(X,R)\to(Y,B)$ be as above satisfying Kawamata’s condition [Ka98] (Definition 4.1). Assume that the horizontal divisor $R_{h}$ is effective, so that its coefficients are in the interval $[0,1)$. Let $\lambda$ be a singular hermitian metric of $L$ given by (a defining meromorphic section of) the divisor $R$.
Then the $L^{2}$ metric $\mu$ for the $\mathbb{Q}$-line bundle $M$ is equal to the product of two singular hermitian metrics $(\mathcal{O}(B_{R}),\eta)$ and $(J:=M-\mathcal{O}(B_{R}),\psi)$, i.e.
$\mu=\eta+\psi\;\;\;\;(e^{-\mu}=e^{-\eta}e^{-\psi})$
where
  • $\eta$ is a singular hermitian metric given by the discriminant divisor $B_{R}$ and
  • $\psi$ is a singular hermitian metric of $J$ with semipositive curvature current and with zero Lelong number at every point of $Y$.
Note that $R$ and $B_{R}$ are not necessarily effective divisors. Thus $\lambda,\mu$ and $\eta$ are not necessarily with semipositive curvature currents. Also unlike $\mu$, the metrics $\psi$ and $\eta$ are not uniquely determined by $\lambda$: they can be added constants $c$ and $-c$, for example.
The discriminant divisor $B_{R}$ defined by [Ka98] is a certain linear combination (see (9)) of the components of $B$, which is supposed to capture the singularities of both $R$ and the singular fibers of $f$, generalizing the case of elliptic fibrations [K63]. The definition of $B_{R}$ is motivated by algebraic geometry : according to the author of [Ka98], the coefficients of the discriminant divisor in [Ka98] were defined so that they behave well under semi-stable reduction.
Our next main result shows that the discriminant divisor has analytic characterization in terms of fiberwise integration along $f$ of singular volume forms. This proves a folklore conjecture, cf. [Ka96, p.81], [Ka98], [Ka99, p.10], [T07]. Together with Theorem 1.1, these answer a question of [EFM18] as well, which asked if there is a metric version of Kawamata’s canonical bundle formula [Ka98].
Theorem 1.2.
Let $f:(X,R)\to(Y,B)$ be as in Theorem 1.1 except that we do not require $R_{h}$ to be effective. Let $m=\dim Y$. Let $\alpha$ and $\beta$ be singular volume forms on $X$ with divisorial poles along the snc $\mathbb{Q}$-divisor $R=\sum a_{i}R_{i}$ and along $R+\lceil-R_{h}\rceil$ respectively, (cf. (3), §2.3). Let $Y_{0}\subset Y$ be the subset of regular values of $f$. Then
  1. (1)
    The fiberwise integration $f_{*}\beta$ of $\beta$ along the smooth fibers $f:X_{0}:=f^{-1}(Y_{0})\to Y_{0}$ has equivalent singularities to a singular volume form $\theta$ (defined on $Y$) with poles along $B_{R}$, the discriminant divisor of $R$, up to a plurisubharmonic weight with vanishing Lelong numbers. In other words, $\theta$ is locally written as
    (1) $\theta(w)=g(w)(\prod^{m}_{i=1}\left|w_{i}\right|^{-2a_{i}})e^{-\psi(w)}\left|dw_{1}\wedge\ldots\wedge dw_{m}\right|^{2}$
    where local analytic coordinates $w=(w_{1},\ldots,w_{m})$ are adapted to $B$ on an open relatively compact subset $U\subset Y$, $\psi$ is a plurisubharmonic function on $U$ with zero Lelong number at every point, $\sum a_{i}\operatorname{div}(w_{i})=B_{R}$ on $U$ and $g:U\to\mathbb{R}$ is a positive and bounded function.
  2. (2)
    The fiberwise integration $f_{*}\alpha$ is less singular than $\theta$.
The statement (1) includes the fiberwise integration version of Theorem 1.1 when $R_{h}\geq 0$ (hence $R+\lceil-R_{h}\rceil=R$). The possible difference of singularities in (2) between $f_{*}\alpha$ and $\theta$ results from components of the snc divisor $R_{h}$ that are not transversal with a special fiber of $f:X_{0}\to Y_{0}$ over a point of $B$, see Remark 5.2. Also note that even if a component $(w_{i}=0)$ of $B$ lies outside $Y_{0}$, the factor $\left|w_{i}\right|^{-2a_{i}}$ appears in (1).
While the fiberwise integration along $f$ is a completely classical operation, Theorem 1.2 is also a metric version in disguise of the canonical bundle formula in terms of the $L^{2}$ metric, which is a metric determined by the fiberwise integration (cf. [PT18, (3.2.2)], [F78]).
1.2. Outline of the proof
$\bullet$ In this paper, we use singular hermitian metrics only for $\mathbb{Q}$-line bundles. First, in the setting of Theorem 1.1, we have the ‘log Calabi-Yau’ condition $K_{X}+L=f^{*}(K_{Y}+M)$ from (11), (12) where $L=O(R)$. When $L$ and $M$ are $\mathbb{Z}$-line bundles, the direct image is $M=f_{*}(K_{X/Y}+L)$ for which the $L^{2}$ metric is defined. The definition and the main properties of the $L^{2}$ metric are generalized when $L$ and $M$ are $\mathbb{Q}$-line bundles in Section 3.
$\bullet$ In the setting of Theorem 1.2, $R_{h}$ is not necessarily effective and we use the direct image $f_{*}(K_{X/Y}+L+A)$ where $A=\mathcal{O}(\lceil-R_{h}\rceil)$. In the $\mathbb{Q}$-line bundle setting, its counterpart is the $\mathbb{Q}$-line bundle $M\otimes(f_{*}A)|_{Y_{1,\text{free}}}$ for which the $L^{2}$ metric is defined, on the Zariski open subset $Y_{1,\text{free}}\subset Y$ where $f_{*}A$ is locally free of rank $1$. (See Theorem 3.7).
$\bullet$ Once the $L^{2}$ metric, say $\mu$, on $Y_{1,\text{free}}$ is prepared as a psh (i.e. plurisubharmonic) metric, our methods of proof of the main theorems in Section 4 proceed via identifying the singularities of the $L^{2}$ metric in terms of all divisorial valuations over $Y$. Since a general psh metric does not have a log resolution, it is necessary to consider all possible divisorial valuations given by prime divisors in proper modifications over $Y$ (cf. [BBJ21, (B.5)]). This is done in the key technical result Theorem 4.5 by carefully applying the defining condition of the discriminant divisor to show that, in fact, the $L^{2}$ metric is valuatively equivalent to the psh metric given by the discriminant divisor.
$\bullet$ This is indeed a very effective viewpoint for our purpose: the psh metric $\psi$ with vanishing Lelong numbers in Theorem 1.1 is obtained rather immediately from comparing the Siu decompositions of the curvature currents of the two metrics in comparison, cf. Lemma 5.1. Namely the divisor part of the Siu decomposition of the curvature current of the $L^{2}$ metric is precisely given by the discriminant divisor, hence the residual part $R_{T}$ itself gives rise to the psh metric $\psi$ for $J$ in the proof of Theorem 1.1 and Theorem 1.2.
$\bullet$ We remark that in Theorem 1.2, it is $f_{*}\beta$, not $f_{*}\alpha$, for which the identification with $\theta$ is given. Note that the choice of adding $T:=\lceil-R_{h}\rceil$ to $R_{h}$ is uniquely determined (among $\mathbb{Q}$-divisors supported on $R_{h}$) so that Theorem 4.5 works: $T$ is the only $\mathbb{Z}$-divisor which makes $T+R_{h}$ both effective and klt. The local integrability from the kltness is indeed used toward the end of the proof of Theorem 4.5.
1.3. Applications to $L^{2}$ extension theorems
Theorem 1.1 and Theorem 1.2 have several diverse applications. We first turn to the application of Theorem 1.2 to $L^{2}$ extension theorems of Ohsawa-Takegoshi type in several complex variables.
Let $Y\subset X$ be a subvariety of a complex manifold. Let $L$ be a holomorphic line bundle on $X$ and $K_{X}$ the canonical line bundle of $X$. An $L^{2}$ extension theorem is a type of a statement that (under suitable conditions on $X,Y,L,\ldots$) if a certain $L^{2}$ norm $\lVert s\rVert_{Y}$ is finite for a holomorphic section $s$ on $Y$ of $(K_{X}\otimes L)|_{Y}$, then there exists $\tilde{s}\in H^{0}(X,K_{X}\otimes L)$ such that $\tilde{s}|_{Y}=s\;\text{ and }\;\lVert\tilde{s}\rVert_{X}\leq c\lVert s\rVert_{Y}$ for some constant $c>0$. Here the $L^{2}$ norm $\lVert s\rVert_{Y}$ plays the crucial role of determining whether $s$ can be extended or not : we will call it as the input norm of the particular statement of $L^{2}$ extension.
Since [OT87], there have been extensive developments on $L^{2}$ extension theorems (cf. e.g. [OT87], [Ma93], [Si96], [D00], [O01], [D15], [K10] among many other works, cf. the introduction of [K23] and references therein.) $L^{2}$ extension theorems are not only analogous to vanishing and injectivity theorems in complex algebraic geometry, but also often serve as their generalizations (cf. [Si02]). In view of the fundamental importance of these powerful tools, it is crucial to understand the criterion of extension given by the above finiteness of the norm $\lVert s\rVert_{Y}$. Such finiteness is nontrivial to achieve due to singularity of the measure against which the norm is taken.
In [D15, Thm. 2.8], Demailly gave an influential $L^{2}$ extension theorem for a log canonical pair $(X,\Psi)$ where $\Psi$ is a quasi-psh function with analytic singularities (see Definition 6.1 for the detailed setting). Let $Y$ be an irreducible component of the non-klt locus of $(X,\Psi)$. The input norm $\lVert s\rVert_{Y}$ is taken with respect to the following Ohsawa measure $dV[\Psi]_{Y}$ which is defined in terms of a certain limit taken over some neighborhoods shrinking to the pole set of $\Psi$.
Definition 1.3 (=Definition 6.1).
[D15, Section 2] Let $dV_{X}$ be a smooth volume form on $X$. The Ohsawa measure $dV[\Psi]_{Y}$ of $\Psi$ on $Y$ (with respect to $dV_{X}$) is a positive measure $d\mu$ on $Y_{\operatorname{reg}}$ satisfying the following condition: for every $g$, a real-valued compactly supported continuous function on $Y_{\operatorname{reg}}$ and for every $\tilde{g}$, a compactly supported extension of $g$ to $X$, we have the relation
[…]
Indeed, in [K23, Thm. 1.2], some essential understanding of the Ohsawa measure was obtained: if $Y$ has non-unique log canonical places, then the Ohsawa measure is the ‘infinity measure’ i.e. only the constant zero function will satisfy the input norm condition. This makes the $L^{2}$ extension result simply unusable on such $Y$, and leaves us the remaining (still very general) case of a unique log canonical place.
In this paper, we complete the geometric investigation of the singularity of the Ohsawa measure by applying Theorem 1.2 when $Y$ has a unique log canonical place. We show that the singularity (or its close upper bound) of the Ohsawa measure $dV[\Psi]_{Y}$ is described by the singularity appearing in Kawamata’s subadjunction from algebraic geometry.
Corollary 1.4 (=see Theorem 6.2).
Let $(X,\Psi)$ and $Y$ be as above, with a unique log canonical place $E\to Y^{\prime}\to Y$ as in (21). The Ohsawa measure $dV[\Psi]_{Y}$ is less singular than the direct image $\pi_{*}v$ under $\pi:Y^{\prime}\to Y$ of a measure $v$ on $Y^{\prime}$ with divisorial poles along Kawamata’s discriminant divisor $B_{R}$ on $Y^{\prime}$ multiplied by a local psh weight with vanishing Lelong numbers.
In particular, this refines the geometric characterization of the $L^{2}$ norm condition given in [K23, Thm. 1.3] with a different method. Corollary 1.4 means that $L^{2}$ extension theorems can indeed extend as many sections as would be expected from algebraic geometry in that, for example when $Y$ does not contain proper log canonical centers in it, $dV[\Psi]_{Y}$ presents no local obstruction for extension as $\pi_{*}v$ does so.
On the other hand, a previous work of the author [K10] first established an $L^{2}$ extension theorem for a log canonical pair, from a different approach not involving the Ohsawa measure. Using Corollary 1.4, we then establish a precise comparison of the input norms of two $L^{2}$ extension theorems : [K10, Thm. 4.2] and [D15, Thm. 2.8] both for a log canonical pair (on a smooth projective variety). Let $\lVert s\rVert_{1}$ be the input norm of [K10, Thm. 4.2] which was taken with respect to a ‘Kawamata metric’, a singularity descibed by the discriminant divisor $B_{R}$ (see Theorem 6.5). Let $\lVert s\rVert_{2}$ be the input norm in [D15, Thm. 2.8] (more precisely in Theorem 6.4) with respect to the Ohsawa measure $dV[\Psi]_{Y}$. We have the following consequence of Corollary 1.4.
Corollary 1.5 (= Theorem 6.5).
For a holomorphic function $s$ on $Y$, if $\lVert s\rVert_{1}$ is locally finite, then $\lVert s\rVert_{2}$ is locally finite.
This means that the combination of Demailly’s $L^{2}$ extension ([D15, Thm. 2.8], [K23, Thm. 3.9]) with the geometric understanding of the Ohsawa measure in Corollary 1.4 amounts to generalization of the $L^{2}$ extension [K10, Thm. 4.2] in that, moreover, a strictly positive curvature condition (“small ample” $A$) in [K10, Thm. 4.2] is replaced with a semipositive one in [D15] (see (23) in Theorem 6.4).
1.4. Applications to semipositivity theorems
As another application of Theorem 1.1, we give a partial answer to a question of Berndtsson (2005) who asked whether it would be possible to use the theory of positivity of direct images (along the line of [B09], [BP08], [PT18], cf. Section 3) in the proof of the celebrated semipositivity theorems in [Ka81], [Ka00] and others, which used Hodge theory.
By establishing Theorem 7.3, we answer this question when the direct image of the relative canonical line bundle is locally free of rank $1$.
Finally, as yet another direction of applications of the relations between the fiberwise integration and the canonical bundle formula as in our main results, we mention recent works in the study of Kähler-Ricci flows, cf. [EGZ18], [LT23].
Remark. Theorem 1.1 in particular recovers nefness of $J$ in [Ka98] by Proposition 2.1. Such refinement of nefness was also given in a recent work [T22] by a different method.  [^2] We mention that the current paper first appeared on arXiv and [T22] was received by a journal, both in October 2019, whereas a manuscript of the current paper was brought to the attention of the author of [T22] in March 2019. We note that the approach of the current paper is not ‘similar’, as opposed to a mention in [T22].
Acknowledgment. The author would like to thank O. Fujino and J. Kollár for answering his questions and having helpful discussions. The author is also grateful to B. Berndtsson, J.-P. Demailly, Y. Kawamata, C. Schnell, V. Tosatti and K.-I. Yoshikawa for helpful conversations and to M. Jonsson, Y. Gongyo and H.S. Kim for useful comments on earlier versions of this work. The author is grateful to S. Boucksom who kindly brought a version of Lemma 4.8 to his attention. This research was partially supported by SRC-GAIA through NRF Korea grant No.2011-0030795 and by Basic Science Research Program through NRF Korea funded by the Ministry of Education (2018R1D1A1B07049683).
2. Singular hermitian metrics and plurisubharmonic functions
We refer to [D11], [B20], [BBJ21, Appendix B] for introduction to singular hermitian metrics of a line bundle and plurisubharmonic (i.e. psh) functions. We have some conventions used in this paper.
  • We will often write holomorphic ($\mathbb{Q}$-)line bundles additively as in $L_{1}+L_{2}:=L_{1}\otimes L_{2}$.
  • We will often denote a singular hermitian metric $e^{-\varphi}$ of a line bundle $L$ simply by $\varphi$. For singular hermitian ($\mathbb{Q}$-)line bundles $(L,\varphi)$, we write additively as in $(L_{1}+L_{2},\varphi_{1}+\varphi_{2})$.
2.1. Singular hermitian metrics on line bundles
A psh metric of a holomorphic $\mathbb{Q}$-line bundle $L$ on a complex manifold $X$ is a singular hermitian metric of $L$ with semipositive curvature current (so that its local weight functions can be taken as psh functions), cf. [D11], [HPS]. When a $\mathbb{Z}$-line bundle $L$ is given transition functions $\{g_{ij}\}$ on a locally trivializing open cover $\{U_{i}\}_{i\in I}$, a (smooth or singular) hermitian metric of $L$ can be identified with a collection of functions $e^{-\varphi_{i}}=\left|g_{ij}\right|^{-2}e^{-\varphi_{j}}$. (A similar description holds when $L$ is a $\mathbb{Q}$-line bundle by taking some multiple $mL$ ($m\geq 1$) that is a genuine holomorphic line bundle.) A holomorphic section $s\in H^{0}(X,L)$ defines a psh metric $\varphi$ of $L$ by taking $\varphi_{i}=\log\left|s_{i}\right|^{2}$ (from $s_{i}=s_{j}g_{ij}$), which can be also denoted by $e^{-\varphi}=\frac{1}{\left|s\right|^{2}}$.
The following fact is due to [D92] (cf. [FF17, (3.5)]). Note that the converse does not hold, see [D11, Example after (6.11)].
Proposition 2.1.
Let $X$ be a compact complex manifold. If $L$ admits a psh metric with vanishing Lelong numbers, then it is nef.
2.2. Valuative equivalence of plurisubharmonic singularities
A function on a complex manifold, with values in ${}\mathbb{R}\cup\{-\infty\}$ is quasi-psh if it is locally the sum of a psh function and a smooth $\mathbb{R}$-valued function. If two quasi-psh functions (or two psh metrics) $\varphi,\psi$ on a complex manifold satisfy that $\varphi-\psi$ is locally bounded above (near every point of the domain), i.e. $\varphi\leq\psi+O(1)$, we say $\psi$ is less singular than $\varphi$, following [D11, (6.3)]. If $\varphi=\psi+O(1)$, we say that $\varphi$ and $\psi$ have equivalent singularities.
As a flexible (and quite different) variant, we will use the following notion (cf. [FJ05, (8.5)] when the dimension is $2$, cf. [BFJ08], [KR22]).
Definition 2.2.
We say that two psh functions $\varphi$ and $\psi$ on a complex manifold $X$ are valuatively equivalent if the following two equivalent conditions (due to [BFJ08] when combined with the strong openness [GZ15]) hold:
(1) For all real $m>0$, the multiplier ideals are equal : $\mathcal{J}(m\varphi)=\mathcal{J}(m\psi)$.
(2) At every point of a proper modification $\pi:X^{\prime}\to X$, the Lelong numbers of $\pi^{*}\varphi$ and $\pi^{*}\psi$ coincide. Equivalently, for every divisorial valuation $v$ centered on $X$, we have $v(\varphi)=v(\psi)$.
When two quasi-psh functions $\varphi$ and $\psi$ on a complex manifold are valuatively equivalent, we note the following facts:
  1. (1)
    If, moreover, $\varphi-\psi$ (or $\psi-\varphi$) happens to be quasi-psh, it should have vanishing Lelong numbers. Of course, in general, none of $\varphi-\psi$ and $\psi-\varphi$ needs to be quasi-psh (even if one of $\varphi,\psi$ has analytic singularities): see, for example, [KS19, (2.3), (2.9)].
  2. (2)
    If $\psi$ has analytic singularities, then $\varphi\leq\psi+O(1)$, cf. [K15, Thm. 4.3].
See e.g. [KS19, (2.8)], [KR22, (4.7)] for more examples and results on valuative equivalence of plurisubharmonic singularities.
2.3. Singular volume forms
A singular volume form $\alpha$ on a complex manifold $X$ can be defined as a general measurable function-valued volume form, i.e. when locally written as $\alpha(w)=g(w)\left|dw_{1}\wedge\ldots\wedge dw_{n}\right|^{2}$ in local analytic coordinates $w=(w_{1},\ldots,w_{n})$, $g\geq 0$ is a local measurable function with values in {$\mathbb{R}\cup\{+\infty\}$. (Here the notation $\left|dw_{1}\wedge\ldots\wedge dw_{n}\right|^{2}$ denotes the standard volume form $dx_{1}\wedge dy_{1}\wedge\ldots\wedge dx_{n}\wedge dy_{n}$ with $dw_{j}=dx_{j}+idy_{j}$ (cf. [BJ17, (1.2)]).)
We will (mostly 333except when $\theta$ in Theorem 1.2 is (locally) extended from $Y_{1}$ to $Y$, which uses the extension property of a psh function) use the case of a continuous singular volume form, namely when the density function $g$ is $[0,\infty]$-valued continuous. A more concrete case for our purpose is as follows.
  • We will say that a singular volume form $\alpha$ on $X$ has (divisorial) poles along an snc $\mathbb{Q}$-divisor (not necessarily effective) $R=\sum^{m}_{i=1}a_{i}R_{i}$ on $X$ if it can be written in local analytic coordinates on $V\subset X$ adapted to $R$ as
    (3) $\alpha(w)=g(w)(\prod^{k}_{i=1}\left|w_{i}\right|^{-2a_{i}})\left|dw_{1}\wedge\ldots\wedge dw_{n}\right|^{2}$
    where $R|_{V}=\sum^{k}_{i=1}a_{i}R_{i}$, $R_{i}=\operatorname{div}(w_{i})$ and $g>0$ is a locally bounded continuous function (i.e. $g$ has no ‘poles or zeros’).
  • Similarly, we will say that a real-valued function $t$ has (divisorial) poles along such $R=\sum a_{i}R_{i}$ if it can be written locally
    (4) $t(w)=g(w)(\prod^{k}_{i=1}\left|w_{i}\right|^{-2a_{i}}).$
For two continuous singular volumes $\alpha$ and $\beta$ on a complex manifold $X$, we will say that $\alpha$ is less singular than $\beta$ if, whenever locally written in terms of local coordinates, $\alpha(w)=g(w)\left|dw_{1}\wedge\ldots\wedge dw_{n}\right|^{2}$, $\beta(w)=h(w)\left|dw_{1}\wedge\ldots\wedge dw_{n}\right|^{2}$ with density functions $g=e^{-u},h=e^{-v}$, we have $u\geq v+O(1)$. (This terminology is compatible with the case when $u$ and $v$ are quasi-psh functions.) If $u=v+O(1)$, we will say that $\alpha$ and $\beta$ have equivalent singularities.
Now let $f:X_{0}\to Y_{0}$ be a proper holomorphic submersion between complex manifolds with $\dim X=m+n$ and $\dim Y=m$. The operation of fiberwise integration sends a singular volume form $\alpha$ on $X$ to a singular volume form $f_{*}\alpha$ on $Y$, whose definition and basic properties are reduced to the local product case by using a partition of unity (cf. [DX, Chapter 1, (2.15) Special case], [N07]). For example, if $\alpha$ is a continuous singular volume form, then so is $f_{*}\alpha$.
3. $L^{2}$ metrics
In [PT18] (cf. [BP08], [HPS]), the authors studied semipositivity of direct images of the form $M:=f_{*}(K_{X/Y}+L)$ for a surjective projective morphism $f:X\to Y$ in terms of certain naturally defined singular hermitian metrics on $M$. We will use the special case of [PT18, (3.2.2)] when $M$ is torsion-free of rank $1$.
3.1. Direct images of adjoint line bundles
In this paper, we use the following special case of [PT18, Thm. 3.3.5] (cf. [HPS, (21.2)]).
Theorem 3.1 (Positivity of direct images, a special case).
[PT18, Thm. 3.3.5], cf. [PT18, Set-up 3.2.1] Let $f:X\to Y$ be a surjective projective morphism with connected fibers between two connected complex manifolds. Let $Y_{0}$ be the set of regular values of $f$.
  • Let $L$ and $M$ be $\mathbb{Z}$-line bundles on $X$ and on $Y$ respectively, such that $K_{X}+L=f^{*}(K_{Y}+M)$ holds.
  • Let $A$ be a $\mathbb{Z}$-line bundle on $X$ such that $f_{*}A$ is torsion-free of rank $1$. Let $Y_{1,\text{free}}$ be the Zariski open subset of $Y$ where $f_{*}A$ is locally free of rank $1$.
  • Let $(L\otimes A,\lambda)$ be a psh metric such that the inclusion
    (5) $f_{*}(K_{X/Y}\otimes L\otimes A\otimes\mathcal{J}(\lambda))\to f_{*}(K_{X/Y}\otimes L\otimes A)$
    is generically an isomorphism.
Then the $L^{2}$ metric $\mu$ for the holomorphic line bundle given by the direct image $M\otimes f_{*}A=f_{*}(K_{X/Y}\otimes L\otimes A)$ restricted to $Y_{0}\cap Y_{1,\text{free}}$ is a psh metric. The $L^{2}$ metric $\mu$ extends to a psh metric for the line bundle given by $M\otimes f_{*}A$ restricted to $Y_{1,\text{free}}$.
We will recall the definition of the $L^{2}$ metric $\mu$ (cf. [PT18, (3.2.2)]) in Definition 3.2. We define the $L^{2}$ metric on $Y_{0}\cap Y_{1,\text{free}}$. For the simplicity of notation, we denote $Y:=Y_{0}\cap Y_{1,\text{free}}$ in the following paragraphs up to Definition 3.2.
Let $u\in H^{0}(Y,M\otimes f_{*}A)$ be a given section. Since $M\otimes f_{*}A=f_{*}(K_{X}+L+A+f^{*}K_{Y}^{-1})$ (where the connected fibers assumption $f_{*}\mathcal{O}_{X}=\mathcal{O}_{Y}$ is used), we have
$H^{0}(Y,M\otimes f_{*}A)=H^{0}(Y,\mathcal{H}om(K_{Y},f_{*}(K_{X}+L+A))).$
Hence viewing $u$ as a sheaf morphism $u:K_{Y}\to f_{*}(K_{X}+L+A)$, for a nowhere vanishing local section $\eta$ of $K_{Y}$ on an open subset $V\subset Y$, we have $u(\eta)\in H^{0}(f^{-1}(V),K_{X}+L+A)$. From [PT18, (3.2.2)], we have locally
(6) $u(\eta)=\sigma_{i}\wedge f^{*}\eta$
for some $\sigma_{i}$, an $(L+A)$-valued holomorphic $n$-form ($n=\dim X-\dim Y$) defined on $U_{i}$ from an open cover $\{U_{i}\}_{i\in I}$ of $f^{-1}(V)$. The existence of such $\sigma_{i}$ follows from computation of elementary nature in terms of local coordinates (‘admissible coordinates’ [MT08, §2.2]) which locally make $f$ a projection. The restrictions $\sigma_{i}|_{X_{y}}$ glue together to define $\sigma|_{X_{y}}$ in the integral below which defines the $L^{2}$ metric $\mu$.
Definition 3.2 ($L^{2}$ metric on $Y_{0}\cap Y_{1,\text{free}}$).
[PT18, (3.2.2)]
Let $(L+A,\lambda)$ be the given hermitian line bundle in Theorem 3.1. We define the $L^{2}$ metric $\mu$ (induced from $\lambda$) for the $\mathbb{Q}$-line bundle $M\otimes f_{*}A$ on $Y:=Y_{0}\cap Y_{1,\text{free}}$ pointwise for $y\in Y$ by letting
(7) $\left(\left|u\right|^{2}\cdot e^{-\mu}\right)(y):=\int_{X_{y}}c_{n}\sigma|_{X_{y}}\wedge\overline{\sigma}|_{X_{y}}e^{-\lambda}=\int_{X_{y}}\left|\sigma_{y}\right|^{2}e^{-\lambda}$
where $u\in H^{0}(Y,M\otimes f_{*}A)$ and $c_{n}=i^{n^{2}}$. Here we write $\sigma_{y}:=\sigma|_{X_{y}}$ noting that $\sigma$ itself may not be globally defined. Also note that the family $\sigma_{y}$ is determined by $u$ only.
We recall the following fiberwise integration property of the $L^{2}$ metric from [PT18], cf. [HPS].
Proposition 3.3.
Let $V\subset Y_{0}$ and $u(\eta)\in H^{0}(f^{-1}(V),K_{X}+L+A)$ be as above. The volume form defined by $\left|u(\eta)\right|^{2}e^{-\lambda}$ on $f^{-1}(V)$ has its fiberwise integration along $f$ equal to $\left|u\cdot\eta\right|^{2}e^{-\mu}$ where $u\cdot\eta\in H^{0}(V,K_{Y}+M)$.
Proof.
As in (6), we have $u(\eta)=\sigma\wedge f^{*}\eta$. Note that this last wedge product makes sense even though $\sigma$ alone may not be defined globally. For simplicity of notation, let us write $f:X\to Y$ in the place of $f:f^{-1}(V)\to V$. Since $\left|u(\eta)\right|^{2}$ is a $\left|L+A\right|^{2}$-valued volume form on $X$444where a section of $\left|L+A\right|^{2}$ is viewed as a collection of local functions compatible with $g_{ij}\overline{g_{ij}}$ for transition functions $g_{ij}$ of $L+A$, its multiplication with a metric $e^{-\lambda}$ for $L+A$ is defined and equal to
$\left|u(\eta)\right|^{2}e^{-\lambda}=\left|\sigma\wedge f^{*}\eta\right|^{2}e^{-\lambda}=\left|\sigma\right|^{2}e^{-\lambda}\cdot\left|f^{*}\eta\right|^{2}$
which is a volume form, i.e. an $(m+n,m+n)$ form on $X$. By the projection formula of fiberwise integration [DX] and using (7), the fiberwise integration of this along $f$ is equal to $\left|u\right|^{2}e^{-\mu}.\left|\eta\right|^{2}.$
The following example illustrates Theorem 3.1.
Example 3.4.
(1) Let $f:X\to Y$ is a ruled surface where $X=\mathbf{P}(E)$ for a holomorphic vector bundle $E$ of rank $2$ on a smooth projective curve $Y$, cf. [H77, §V.2]. When $L=O_{\mathbf{P}(E)}(2)$ and $M=\det E$, we have $K_{X}+L=f^{*}(K_{Y}+M)$. When $L$ is ample and provided with a smooth psh metric, Theorem 3.1 says that $M$ is semipositive, which is consistent with [H77, §V, (2.20), (2.21)], characterization of ample line bundles on $X$.
(2) Now let $X=\mathbf{P}(E)$ be as in [DPS94, (1.7)] with $L=O_{\mathbf{P}(E)}(2)$ : it is shown there that the only possible psh metrics $\lambda$ for $L$ are ones whose curvature current is equal to $2[C]$ where $C\subset X$ is a section of $f$ (in particular, $L$ is not semipositive). It is clear that this $\lambda$ does not satisfy the condition (5) of Theorem 3.1.
3.2. Generalization to $\mathbb{Q}$-line bundles
In this subsection, we will generalize Theorem 3.1 to Theorem 3.7. Let $L$ and $M$ be $\mathbb{Q}$-line bundles satisfying the relation $K_{X}+L=f^{*}(K_{Y}+M)$. Note that even when $K_{X},L$ and $K_{Y}$ are $\mathbb{Z}$-line bundles, it is possible that $M$ is only a $\mathbb{Q}$-line bundle (see e.g. (27)).
Recall that the only case of the direct images we use from [PT18, Thm. 3.3.5] is of the form $M\otimes f_{*}A$ in Theorem 3.1, and we only use its restriction to $Y_{1,\text{free}}$ which is a line bundle. In the case of a $\mathbb{Q}$-line bundle $M$, we will use the notation $M\otimes f_{*}A$ only to denote the $\mathbb{Q}$-line bundle $M\otimes(f_{*}A|_{Y_{1,\text{free}}})$ on $Y_{1,\text{free}}$. We will also refer to it as the ‘$\mathbb{Q}$-line bundle given by $M\otimes f_{*}A$ restricted to $Y_{1,\text{free}}$’. The $L^{2}$ metric will be defined as a singular hermitian metric for this $\mathbb{Q}$-line bundle, as follows.
Let $N=(k-1)M$ where $k\geq 1$ is the smallest integer such that $kM$ is a $\mathbb{Z}$-line bundle (i.e. a usual holomorphic line bundle). Then $K_{X}+L+f^{*}N=f^{*}(K_{Y}+M+N)$ is a $\mathbb{Z}$-line bundle on $X$.
Definition 3.5.
Let $(L+A,\lambda)$ be a psh metric. Let $(N,\chi)$ be a smooth hermitian metric (without requiring curvature condition), which plays an auxiliary role. Equip $L+A+f^{*}N$ with the metric $\lambda+f^{*}\chi$. Applying Definition 3.2 of the $L^{2}$ metric in the $\mathbb{Z}$-line bundle case, we get the $L^{2}$ metric $\mu_{\chi}$ for $(M+N)\otimes f_{*}A$, induced by $\lambda+f^{*}\chi$. Multiplying $(-N,-\chi)$, we get a metric $\mu_{\chi}-\chi$ for the $\mathbb{Q}$-line bundle $M\otimes f_{*}A$ (restricted to $Y_{0}\cap Y_{1,\text{free}}$). We define $\mu_{\chi}-\chi$ to be the $L^{2}$ metric induced by the given $(L+A,\lambda)$, which is justified by the following.
Proposition 3.6.
The $L^{2}$ metric $\mu:=\mu_{\chi}-\chi$ for the $\mathbb{Q}$-line bundle $M\otimes f_{*}A$ (restricted to $Y_{0}\cap Y_{1,\text{free}}$) is well-defined, independent of the choice of $\chi$.
Proof.
We need to check that $\mu_{\chi}-\chi$ is independent of the choice of the smooth metric $\chi$. It suffices to check this at each point $y\in Y_{0}\cap Y_{1,\text{free}}$, hence we may fix a local trivialization of the holomorphic line bundle $kM$ in a neighborhood $V$ of $y$. Then $\mu_{\chi}$ and $\chi$ are functions on $V$. Now let $u$ be a holomorphic section of $M+N=kM$. From (7), we have
(8) $\left|u\right|^{2}e^{-{\mu}_{\chi}}(y)=\int_{X_{y}}\left|\sigma_{y}\right|^{2}e^{-\lambda}e^{-f^{*}\chi}=e^{-\chi(y)}\int_{X_{y}}\left|\sigma_{y}\right|^{2}e^{-\lambda}$
where $\sigma_{y}$ is as in Definition 3.2. We get the second equality since $f^{*}\chi$ is constant on $X_{y}$. ∎
Also using this definition involving $M+N=kM$ and $\chi$, we have the fiberwise integration property for the $L^{2}$ metric inherited from Proposition 3.3.
Now we have the following version of Theorem 3.1 when $L$ and $M$ are $\mathbb{Q}$-line bundles.
Theorem 3.7.
Let $f:X\to Y$ be a surjective projective morphism with connected fibers between two connected complex manifolds. Let $Y_{0}$ be the set of regular values of $f$.
  1. (1)
    Let $L$ and $M$ be $\mathbb{Q}$-line bundles on $X$ and on $Y$ respectively, such that $K_{X}+L=f^{*}(K_{Y}+M)$ holds as an equality of $\mathbb{Q}$-line bundles.
  2. (2)
    Let $A$ be a $\mathbb{Z}$-line bundle on $X$ such that $f_{*}A$ is torsion-free of rank $1$. Let $Y_{1,\text{free}}$ be the Zariski open subset of $Y$ where $f_{*}A$ is locally free of rank $1$.
  3. (3)
    Let $(L+A,\lambda)$ be a psh metric such that $\mathcal{J}(\lambda|_{F})=\mathcal{O}_{F}$ for a general fiber $F$ of $f$.
Then the $L^{2}$ metric $\mu$ for the $\mathbb{Q}$-line bundle given by $M\otimes f_{*}A$ defined on $Y_{0}\cap Y_{1,\text{free}}$ is a psh metric and it extends to a psh metric for the $\mathbb{Q}$-line bundle given by $M\otimes f_{*}A$ on $Y_{1,\text{free}}$.
The case when $A=\mathcal{O}_{X}$ will be used in the proof of Theorem 1.1 while the general case will be used in the proof of Theorem 1.2.
Proof.
Both assertions (being a psh metric and extension) are local properties. As before, let $N=(k-1)M$ where $k\geq 1$ is the smallest integer such that $kM$ is a $\mathbb{Z}$-line bundle. For $y\in Y_{0}\cap Y_{1,\text{free}}$, fix a local trivialization of $kM$ in a neighborhood $V$ of $y$. Take $\chi=0$ and apply Theorem 3.1 to $f^{-1}(V)\to V$ to obtain the psh property. Similarly, extension to $Y_{1,\text{free}}$ follows from applying Theorem 3.1 to $f^{-1}(V)\to V$ where $V$ is a neighborhood of points $y\in Y_{1,\text{free}}\setminus Y_{0}$. ∎
4. The setting and the valuative equivalence result
In this section, we give the proof of the main theorems.
4.1. The setting of Kawamata’s canonical bundle formula
We first give the setting of the main theorems, which is the same as in [Ka98] (cf. [Ko07]) except that we are in the generality of complex manifolds (where the necessary Hironaka resolution theorems also hold, cf. [H64], [AHV77], [W09]).
Let $(X,R)$ and $(Y,B)$ be two pairs of complex manifolds and snc $\mathbb{Q}$-divisors. Let $f:X\to Y$ be a surjective projective morphism with connected fibers. An irreducible component $R_{i}$ of $R$ is called horizontal if $f(R_{i})=Y$. Otherwise it is called vertical. We write $R=R_{h}+R_{v}$ where $R_{h}$ is the horizontal part and $R_{v}$ is the vertical part. For a divisor $R=\sum a_{i}R_{i}$ ($a_{i}\neq 0$), define $\operatorname{red}(R):=\sum R_{i}$.
Definition 4.1.
[Ko07, Def. 8.3.6], [Ka98, Thm. 2] We will say that $f:(X,R)\to(Y,B)$ satisfies Kawamata’s condition if the following hold:
  1. (1)
    $X$, $Y$ are complex manifolds.
  2. (2)
    $B$ is a reduced snc divisor on $Y$, i.e. $B=\operatorname{red}(B)$.
  3. (3)
    $f(\operatorname{Supp}(R_{v}))\subset B$.
  4. (4)
    $\operatorname{red}(R)+f^{*}B$ is an snc divisor on $X$.
  5. (5)
    $f$ is a submersion over $Y\setminus B$.
  6. (6)
    $R_{h}$ is relative snc over $Y\setminus B$.
  7. (7)
    Coefficients of $R_{h}$ are in the interval $(-\infty,1)$.
  8. (8)
    $\operatorname{rank}f_{*}\mathcal{O}_{X}(\lceil-R_{h}\rceil)=1$.
  9. (9)
    $K_{X}+R$ is $\mathbb{Q}$-linearly equivalent to the pullback of some $\mathbb{Q}$-Cartier divisor on $Y$.
Let $B_{R}$ be the discriminant divisor induced by $R$ which is defined as (when $B=\sum B_{i}$) $B_{R}=\sum c_{i}B_{i}$ where
(9) $c_{i}:=1-\sup\{c:(X,R+cf^{*}B_{i})\text{\; is log canonical over the general point of }B_{i}\}$
([Am99, Def. 3.1]). The condition in the sup is that, in other words, the non-lc locus of the snc divisor $R+cf^{*}B_{i}$ is contained in the inverse image of the union of some Zariski closed proper subsets of $B_{i}$’s. Since $\operatorname{red}(R)+f^{*}B$ is snc, the lc (resp. klt) conditions are determined by coefficients of all the components being $\leq 1$ (resp. $<1$), cf. [Ko97, Cor. 3.12].
Now write $R_{v}=R_{v1}+R_{v2}$ where the sum $R_{v1}$ (resp. $R_{v2}$) consists of components whose images under $f$ are of codimension $1$ (resp. of codimension at least $2$) in $Y$. Since the coefficients of $R_{h}$ are less than $1$, the condition in (9) is then equivalent to $(X,R_{v1}+cf^{*}B_{i})\text{\; being lc over the general point of }B_{i}$. Hence $B_{R}$ is also characterized as the unique smallest $\mathbb{Q}$-divisor supported on $B$ such that
(10) $R_{v1}+f^{*}(B-B_{R})\leq\operatorname{red}(f^{*}B)$
cf. [Ko07, (8.3.7) (2)].  [^5] We thank Hyunsuk Kim for pointing out to replace $R_{v}$ by $R_{v1}$ in (10). (Also see [Ka98, p.895] for the equivalent original definition of $B_{R}$.)
Remark 4.2.
It can be easily seen (from local expression of $f$ in terms of coordinates adapted to the snc divisors) that the components of $R_{v1}$ are contained in the components of $f^{*}B$.
Later we will need the following lemma.
Lemma 4.3.
In the setting of (4.1), let $S$ be a $\mathbb{Q}$-divisor supported on $B$. The discriminant divisor $B_{R+f^{*}S}$ associated to $R+f^{*}S$ is equal to $B_{R}+S$.
Proof.
It follows from the definition putting $R^{\prime}=R+f^{*}S$ and $R^{\prime}_{v}=R_{v}+f^{*}S$. ∎
In the setting of Definition 4.1 (simply using the condition (8)), a $\mathbb{Q}$-line bundle $J(X/Y,R)$ (called the moduli part) is defined on $Y$ by the following relation (cf. [Ka98])
(11) $K_{X}+R\sim_{\mathbb{Q}}f^{*}(K_{Y}+B_{R}+J(X/Y,R)).$
For the purpose of this paper involving metrics, we also view (11) as the corresponding equality of $\mathbb{Q}$-line bundles
(12) $K_{X}+L=f^{*}(K_{Y}+M)=f^{*}(K_{Y}+J+H)$
where we define the corresponding $\mathbb{Q}$ line bundles $L:=\mathcal{O}(R),H:=\mathcal{O}(B_{R}),\;J:=J(X/Y,R)$ and $M:=J+H$.
Also we record the following obvious statement for later use.
Lemma 4.4.
Let $R_{h}=\sum a_{i}D_{i}$ be the snc divisor in Definition 4.1 with the coefficients $a_{i}$ in the interval $(-\infty,1)$. Then the divisor $R_{h}+\lceil-R_{h}\rceil$ has all the coefficients in the interval $[0,1)$.
Let $A=\mathcal{O}(\lceil-R_{h}\rceil)$ be the corresponding $\mathbb{Z}$-line bundle on $X$. Note that $A=\mathcal{O}_{X}$ in the setting of Theorem 1.1.
4.2. Valuative equivalence of the $L^{2}$ metric
In the setting of the previous subsection, the following is the key technical result in the proof of the main results.
Theorem 4.5.
Let $f:X\to Y$, $K_{X}+L=f^{*}(K_{Y}+M)$ and $A=\mathcal{O}(\lceil-R_{h}\rceil)$ be as in the previous subsection. Assume that the snc divisor $R_{v1}$ (i.e. those vertical components whose images are of codimension $1$ in $Y$) is effective. (Hence the discriminant divisor $B_{R}$ is effective.)
Let $\lambda$ be a singular hermitian metric of $L+A$ given by the divisor $R+\lceil-R_{h}\rceil$ (so that $\mathcal{J}(\lambda|_{F})=\mathcal{O}_{F}$ holds for a general fiber $F$ of $f$ due to Definition 4.1 (6), (7)). Let $\mu$ be the $L^{2}$ metric (induced from $\lambda$) for $M\otimes f_{*}A$, defined in Proposition 3.6. Then the following hold.
  1. (1)
    The $L^{2}$ metric $\mu$ is a psh metric for the $\mathbb{Q}$-line bundle given by $M\otimes f_{*}A$ restricted to the Zariski open subset $Y_{1,\text{free}}\subset Y$ where $f_{*}A$ is locally free (of rank $1$).
  2. (2)
    Let $\varphi_{B_{R}}$ be a psh metric for the $\mathbb{Q}$-line bundle $\mathcal{O}(B_{R})$ given by the $\mathbb{Q}$-divisor $B_{R}\geq 0$. The $L^{2}$ metric $\mu$ is valuatively equivalent to $\varphi_{B_{R}}$ on $Y_{1,\text{free}}$.
Proof.
(1) Let $Y^{\prime}:=Y\setminus Z$ where $Z$ is the image of the components in $R_{v2}$ (hence $Z$ is of codimension $\geq 2$). Apply Theorem 3.7 to the restriction $f^{-1}(Y^{\prime})\to Y^{\prime}$ noting that the condition (2) of Theorem 3.7 is satisfied for $A=\mathcal{O}(\lceil-R_{h}\rceil)$, cf. [Ka98], [Am04].
Note that $\lambda$ given by (i.e. having divisorial poles along) the divisor $R+\lceil-R_{h}\rceil=R_{h}+\lceil-R_{h}\rceil+R_{v1}+R_{v2}$ is a psh metric on $f^{-1}(Y^{\prime})$ since it is taking the complement of the possibly noneffective divisor $R_{v2}$. Hence the $L^{2}$ metric $\mu$ is a psh metric on $Y_{1,\text{free}}\setminus Z$ by Theorem 3.7. Since $Z$ is of codimension $\geq 2$, $\mu$ extends to $Y_{1,\text{free}}$ as a psh metric of the same $\mathbb{Q}$-line bundle (as its psh local weight functions extend so).
(2) Let $\tilde{R}:=R+\lceil-R_{h}\rceil$. Note that, from the definition, $B_{\tilde{R}}=B_{R}$. In the rest of this proof, for the convenience of notation, we will use $R$ to denote $\tilde{R}$.
We need to show that $v(\mu)=v(B_{R})(:=v(\varphi_{B_{R}}))$ for every divisorial valuation $v=\operatorname{ord}_{G}$ where $G$ is a prime divisor over $Y$ with nonempty center in $Y$, cf. [BBJ21, Section B.5]. More precisely, there exists a proper modification $\pi:Y^{\prime}\to Y$ such that $G\subset Y^{\prime}$ is a prime divisor and its center on $Y$ is $\pi(G)\neq\emptyset$.
Now we consider the following diagram (13) where $\rho:X^{\prime}\to X$ is bimeromorphic and $f^{\prime}$ also satisfies Kawamata’s condition (4.1). (The morphism $f^{\prime}:X^{\prime}\to Y^{\prime}$ in this diagram is obtained by first taking the fibration $X^{\prime\prime}\to Y^{\prime}$ induced by the bimeromorphic base change $Y^{\prime}\to Y$ and then taking further blow-ups over $X^{\prime\prime}$ : cf. [Ko07, 8.4.8].)
We take a Zariski open subset $Y_{0}\subset Y$ and let $X_{0}:=f^{-1}(Y_{0})$ so that $f:X_{0}\to Y_{0}$ is submersion. Also determine $Y^{\prime}_{0}$ and $X^{\prime}_{0}:=f^{\prime-1}(Y^{\prime}_{0})$ in the following diagram so that they are isomorphic to $Y_{0}$ under $\pi$ and to $X_{0}$ under $\rho$ respectively.
[…]
The restrictions $f:X_{0}\to Y_{0}$ and $f^{\prime}:X^{\prime}_{0}\to Y^{\prime}_{0}$ can be identified with each other and so are the fiberwise integrations taken along them. This can be expressed as
(14) $\pi^{*}(f_{*}u)=f^{\prime}_{*}(\rho^{*}u)$
for a singular volume form $u$, where the domains are all restricted to $X_{0},Y_{0},X^{\prime}_{0},Y^{\prime}_{0}$. Now let $R^{\prime}$ be the divisor on $X^{\prime}$ defined by $K_{X^{\prime}}+R^{\prime}=\rho^{*}(K_{X}+R)$. Now we take $u$ to be a singular volume form with poles along $R$, so that the pullback $\rho^{*}u$ has poles along $R^{\prime}=\rho^{*}R-K_{X^{\prime}/X}$.
Consider the following projection formula of fiberwise integration combined with (14):
(15) $f^{\prime}_{*}(\rho^{*}u\wedge f^{\prime*}t)=f^{\prime}_{*}(\rho^{*}u)\wedge t=\pi^{*}(f_{*}u)\wedge t.$
In order to show $v(\mu)=v(B_{R})$, we will apply (15) for the fiberwise integration taken along the restriction of $f^{\prime}$ over a neighborhood $U\subset Y^{\prime}$ of a general point of $G$ taking $t$ with poles $\delta B^{\prime}-(v(\mu)-\alpha)G$ first and then with poles $\delta B^{\prime}-(v({B_{R}})-\alpha)G$ secondly. Here $\alpha:=\operatorname{ord}_{G}(K_{Y^{\prime}/Y})$, i.e. the coefficient of the prime divisor $G$ in the relative canonical divisor $K_{Y^{\prime}/Y}$.
For arbitrary $\delta<1$, take a function $t$ with poles $\delta B^{\prime}-(v(\mu)-\alpha)G$. We claim that
(16) $R^{\prime}_{v1}+f^{\prime*}(\delta B^{\prime}-(v(\mu)-\alpha)G)<\operatorname{red}(f^{\prime*}B^{\prime}).$
To verify the claim (16), first note that since $f_{*}u$ has the poles given by the psh weight $e^{-\mu}$, its pullback $\pi^{*}(f_{*}u)$ on $Y^{\prime}$ has poles given by the pullback of $e^{-\mu}$ divided by the contribution of the jacobian of the morphism $Y^{\prime}\to Y$. When $g=0$ is a local equation of the divisor $G$, this pole is expressed as $e^{-\pi^{*}\mu}\left|g\right|^{2\alpha}$.
Since we are looking at a neighborhood $U$ of a general point of $G$ and considering the restriction of the divisor $\delta B^{\prime}-(v(\mu)-\alpha)G$ to $U$, which is nothing but $(\delta-v(\mu)+\alpha)G$, the local expression of the pole of $\pi^{*}(f_{*}u)\wedge t$ in (15) is now given by
$e^{-\pi^{*}\mu}\left|g\right|^{2\alpha}\left|g\right|^{-2(\delta-v(\mu)+\alpha)}=e^{-\pi^{*}\mu}\left|g\right|^{2(v(\mu)-\delta)}.$
By Lemma 4.8 applied for $W=Y^{\prime},\varphi=\pi^{*}\mu,\psi=2(v(\mu)-\delta)\log\left|g\right|$, this is locally integrable. Thus from the LHS of (15), the singular volume form inside $f^{\prime}_{*}$ is also locally integrable, which gives (16).
Now from (16), it follows that $R^{\prime}_{v1}+f^{\prime*}(B^{\prime}-(v(\mu)-\alpha)G)\leq\operatorname{red}(f^{\prime*}B^{\prime})$. From the definition of the discriminant divisor $B^{\prime}_{R^{\prime}}$ associated to $R^{\prime}$ (see (10)), we see that $v(B^{\prime}_{R^{\prime}})$ is the smallest possible coefficient for $G$ (in the place of ($v(\mu)-\alpha$)) to make this inequality hold. Hence we have
(17) $v(\mu)-\alpha\geq v(B^{\prime}_{R^{\prime}})=v(B_{R})-\alpha$
where the equality is from Lemma 4.7. Thus we have $v(\mu)\geq v(B_{R})$.
Now suppose that $v(\mu)-v(B_{R})>0$. This time, take a function $t$ to be with poles $\delta B^{\prime}-(v(B_{R})-\alpha)G$ for $\delta<1$. From (15), we will have contradiction since the LHS is locally integrable while the RHS is not. Indeed, the RHS of (15) has poles along
$(v(\mu)-\alpha)G+\delta B^{\prime}-(v(B_{R})-\alpha)G=(v(\mu)-v(B_{R}))G+\delta B^{\prime}$
and it is not locally integrable when $v(\mu)-v(B_{R})+\delta>1$ (note that $G$ appears in $B^{\prime}$ as we may assume so).
On the other hand, the LHS of (15) is locally integrable since $\rho^{*}u\wedge f^{\prime*}w$ is locally integrable before taking the fiberwise integration : it has poles along the snc divisor $T:=R^{\prime}+f^{\prime*}(\delta B^{\prime}-(v(B^{\prime}_{R^{\prime}})G)$ which is klt (since $\delta<1$) for the following reason. Since $T$ is snc, we only need to check the coefficients of the components. For the vertical part of $T$ (with respect to $f^{\prime}$), this follows from the definition of the discriminant divisor $B^{\prime}_{R^{\prime}}$.
Each horizontal component of $T$ is contained in $R^{\prime}$ and is the strict transform under $\rho:X^{\prime}\to X$ of a horizontal component for $f$ from the construction of $X^{\prime}$. In view of $K_{X^{\prime}}+R^{\prime}-\rho^{*}R_{v}=\rho^{*}(K_{X}+R_{h})$, the coefficients of the horizontal part of $T$ is also less than $1$. Note that our notation $R$ really means $\tilde{R}=R+\lceil-R_{h}\rceil$ (for the given $R$ in the statement), where $\tilde{R}_{h}=R_{h}+\lceil-R_{h}\rceil$ is klt due to Lemma 4.4.
This is contradiction. Hence we have $v(\mu)=v(B_{R})$, which completes the proof of Theorem 4.5. ∎
Remark 4.6.
At the end of this argument, we see that adding precisely $S:=\lceil-R_{h}\rceil$ to $R_{h}$ is essential for this result and its proof : $S$ must be an integral divisor (so that we can take the line bundle $A=\mathcal{O}(S)$) and $S+R_{h}$ should be effective.
Lemma 4.7.
For $f$ and $f^{\prime}$ in (13) satisfying Kawamata’s condition, we have $v(B^{\prime}_{R^{\prime}})=v(B_{R})-\alpha$.
Proof.
This follows from the equality of divisors $K_{Y^{\prime}}+B^{\prime}_{R^{\prime}}=\pi^{*}(K_{Y}+B_{R})$ which in turn follows immediately from the pull back property of the moduli part line bundles $J(X^{\prime}/Y^{\prime},R^{\prime})=\pi^{*}J(X/Y,R)$ [Ko07, 8.4.9 (3)] (originally due to [Ka98]). The pull back property results from the Hodge theoretic characterization of the moduli part line bundle and the fact that the canonical extension in Hodge theory commutes with pull backs. ∎
Also the following lemma was used, a weaker variant (using only one divisorial valuation) of valuative characterization of multiplier ideals (cf. [BFJ08], [BBJ21]).
Lemma 4.8.
Let $W$ be a complex manifold and $H$ a prime divisor on $W$. Let $\varphi$ and $\psi$ be psh functions on $W$. Assume that $\psi$ has analytic singularities. If $\operatorname{ord}_{H}(\psi)>\operatorname{ord}_{H}(\varphi)-1$, then $e^{\psi-\varphi}$ is locally integrable at a general point of $H$.
666 Note the minor point that here $e^{\psi-\varphi}$ is locally $L^{1}$ instead of $L^{2}$ as in [B20, Thm. 10.11] since our convention from §2.1 is taking, for example, $\varphi=\log\left|f\right|^{2}$ instead of $\varphi=\log\left|f\right|$ throughout this paper.
Proof.
(1) First assume that $\varphi$ also has analytic singularities. Let $\mu:W^{\prime}\to W$ be a log resolution of both $\varphi$ and $\psi$. We may choose a general point $p$ of $H$ and shrink $W$ to some $U$, a neighborhood of $p$, in order to avoid prime divisors on $W$ along which $\varphi$ or $\psi$ has positive generic Lelong numbers and also to avoid the images of exceptional divisors of $\mu$ along which $\mu^{*}\varphi$ or $\mu^{*}\psi$ has positive generic Lelong numbers. Then it is easy to see from $\mu:\mu^{-1}(U)\to U$ that the condition $v(\psi-\varphi)>-1$ for $v=\operatorname{ord}_{H}$ alone determines local integrability of $e^{\psi-\varphi}$.
(2) Now let $\varphi$ be a general psh function. Let $(\varphi_{m})_{m\geq 1}$ be a Demailly approximation sequence of $\varphi$. (cf. [D13, Thm. 7] for the case of a bounded pseudoconvex domain in $\mathbb{C}^{n}$. In general, one can use partitions of unity as in [D11] to choose such a sequence.) Since $v(\psi)>v(\varphi)-1$, we can fix $\lambda>1$ such that $v(\psi)>\lambda v(\varphi)-1$. Then for every $m\geq 1$, we have $v(\psi)>\lambda v(\varphi_{m})-1$ since $\varphi_{m}$ is less singular than $\varphi$ in the sense that $\varphi_{m}\geq\varphi+O(1)$. Applying (1) to $\varphi_{m}$ (having analytic singularities), we then have $e^{\psi-\lambda\varphi_{m}}\in L^{1}_{\operatorname{loc},x}$ at a general point $x$ of $H$, i.e. for $x\in H\setminus V_{m}$ where $V_{m}\subset H$ is a Zariski closed subset (which may depend on $m$). By Lemma 4.9, we have $e^{\psi-\varphi}\in L^{1}_{\operatorname{loc},x}$ for the same $x$. ∎
Lemma 4.9.
[D13] Let $\psi,\varphi$ be psh functions on a complex manifold $X$. Let $(\varphi_{m})_{m\geq 1}$ be a Demailly approximation sequence of $\varphi$. Let $\lambda>1$ be a real number. For every $x\in X$ and every integer $m\geq\lceil\frac{\lambda}{2(\lambda-1)}\rceil$, we have the implication
$e^{\psi-\lambda\varphi_{m}}\in L^{1}_{\operatorname{loc},x}\implies e^{\psi-\varphi}\in L^{1}_{\operatorname{loc},x}.$
Proof.
When $\psi=\log\left|f\right|^{2}$ for a holomorphic function $f$, this is [D13, Cor. 4] which states the inclusion of multiplier ideals $\mathcal{J}(\lambda\varphi_{m})\subset\mathcal{J}(\varphi)$. The same argument (which uses [D13, Lem. 2]) works when we put $e^{\psi}$ in the place of $\left|f\right|^{2}$. ∎
5. Proofs of the main theorems
Now using Theorem 4.5, we complete the proof of Theorem 1.1.
Proof of Theorem 1.1.
First note that $R_{v}$ is not necessarily effective. By Remark 4.2, there exists an effective divisor $S$ supported on $B$ such that $R_{v1}+f^{*}S\geq 0$ and $B_{R}+S\geq 0$. Let $N=\mathcal{O}(S)$ be the associated $\mathbb{Q}$-line bundle. Consider the equality of $\mathbb{Q}$-line bundles
$K_{X}+L+f^{*}N=f^{*}(K_{Y}+J+H+N).$
Equip $L+f^{*}N$ with a (not necessarily psh) singular hermitian metric $\lambda$ given by (i.e. having divisorial poles along) the divisor $R+f^{*}S$. Then the induced $L^{2}$ metric $\mu$ for $J+H+N$ is a psh metric by Theorem 4.5 (1), which we are applying to the snc divisor $R+f^{*}S=R_{h}+R_{v1}+R_{v2}+f^{*}S$ in the place of ‘$R$ in the statement of Theorem 4.5’ in view of $R_{v1}+f^{*}S\geq 0$. Here note that the condition $\mathcal{J}(\lambda|_{F})=\mathcal{O}_{F}$ for a general fiber $F$ in Theorem 3.7 is satisfied since the coefficients of the horizontal snc divisor $R_{h}$ are assumed to be less than $1$.
Now consider the Siu decomposition [Si74], cf. [D11, (2.18)], [B04, 2.2.1] of the curvature current $\Theta_{\mu}$ of $\mu$:
$T:=\Theta_{\mu}=\sum_{W}\nu(T,W)[W]+R_{T}$
where $W$ runs through all codimension $1$ irreducible analytic subsets of $X$ and $\nu(T,W)$ is the generic Lelong number of $T$ along $W$. The closed semipositive current $R_{T}$ is the residual part of the decomposition whose Lelong superlevel sets $E_{c}(R_{T})$ have codimension $\geq 2$ for every $c>0$.
By Theorem 4.5 (2), the current $\Theta_{\mu}$ (or its psh potential) is valuatively equivalent to the current given by the effective discriminant divisor $B_{R+f^{*}S}=B_{R}+S\geq 0$ (by Lemma 4.3). Thus the divisor part $\sum\nu(T,Y_{k})[Y_{k}]$ is a finite sum which is precisely given by the discriminant divisor $B_{R+f^{*}S}$ of the snc divisor $R+f^{*}S$.
We apply Lemma 5.1 to the curvature current $T$ of $(J+H+N,\mu)$ and the curvature current $Q$ of $(H+N,\varphi_{B_{R}+S})$ where $\varphi_{B_{R}+S}$ is a psh metric given by the effective divisor $B_{R}+S$. Since the closed positive $(1,1)$ current $R_{Q}=0$, the current $R_{T}$ has zero Lelong numbers at every point by Lemma 5.1.
Since the closed positive $(1,1)$ current $R_{T}$ belongs to the first Chern class of the $\mathbb{Q}$-line bundle $J+H+N-(H+N)=J$, there exists a singular hermitian metric $\psi$ of $J$ whose curvature current is equal to $R_{T}$ (as is well-known, see e.g. [B04, p.50]). This $\psi$ is the psh metric we wanted in the statement of (1) of the theorem : it has vanishing Lelong numbers.
Choose a singular hermitian metric $\varphi_{S}$ given by the divisor $S$ such that $\varphi_{B_{R}+S}=\varphi_{B_{R}}+\varphi_{S}$. From $(J+H+N,\mu)$, we subtract $(J,\psi)$ and get a psh metric $(H+N,\mu-\psi)$ given by the effective divisor $B_{R}+S$. Now subtracting again $(N,\varphi_{S})$, we get $(H,\eta:=\mu-\psi-\varphi_{S})$ which is a singular hermitian metric given by the original discriminant divisor $B_{R}$. Since $B_{R}$ may not be effective, $\eta$ may not be a psh metric. This $\eta$ is the one we were looking for, which completes the proof of Theorem 1.1. ∎
Lemma 5.1.
Let $X$ be a complex manifold. Let $T$ and $Q$ be closed semipositive $(1,1)$ currents on $X$. Suppose that $T$ and $Q$ are valuatively equivalent, i.e. they have the same Lelong numbers at every point in $X$ and at every point in all proper modifications $\tilde{X}\to X$. Then in the Siu decomposition of the closed positive $(1,1)$ currents $T$ and $Q$,
$T=\sum_{W}\nu(T,W)[W]+R_{T}\;\;\;\text{and}\;\;\;Q=\sum_{W}\nu(Q,W)[W]+R_{Q},$
the two residual parts, i.e. the closed semipositive $(1,1)$ currents $R_{T}$ and $R_{Q}$ are valuatively equivalent.
Proof.
This is immediate from the construction of $R_{T}$ (and $R_{Q}$) of the Siu decomposition (cf. [B04, 2.2.1]) as the limit $R_{T}=T-\sum_{k\geq 1}\nu(T,W_{k})[W_{k}]$ where $(W_{k})_{k\geq 1}$ is an at most countable family of irreducible analytic subsets of codimension $1$. Note that we have $\nu(T,W)=\nu(Q,W)$ for every $W$. Also note that $v(R_{T})=v(R_{Q})$ is nonzero only possibly for $v=\operatorname{ord}_{G}$ where $G$ is a prime divisor over $X$ (i.e. $G\subset X^{\prime}\to X$ for a proper modification $\pi:X^{\prime}\to X$) whose center on $X$, $\pi(G)$ is of codimension $\geq 2$ in $X$. ∎
Now we turn to the proof of Theorem 1.2.
Proof of Theorem 1.2.
We will use the full setting of Theorem 4.5 to the given $f:(X,R)\to(Y,B)$, taking $A=\mathcal{O}(\lceil-R_{h}\rceil)$. Let $Y_{1}:=Y_{1,\text{free}}$ be the Zariski open subset of $Y$ where $f_{*}A$ is locally free of rank $1$. The result will follow from a version of Theorem 1.1 for the morphism $f:f^{-1}(Y_{1})\to Y_{1}$.
Step 1. Since $\alpha$ is a continuous singular volume form, so is its fiberwise integration $f_{*}\alpha$ on $Y_{0}$. Hence we may check ‘less singular’ on $Y_{0}\cap Y_{1}$ once $\theta$ is constructed from the $L^{2}$ metric in this case, which will be a singular hermitian metric defined on $Y_{1}$.
Since being ‘less singular’ is a local property on $Y_{0}\cap Y_{1}$, we may restrict to an open neighborhood of a point $p\in Y_{0}\cap Y_{1}$. (The $\theta$ we consider will extend from $Y_{1}$ to $Y$, then we use a partition of unity to obtain $\theta$ on $Y$ as in the statement.)
Step 2. Now we take the following variant of Theorem 1.1 for $f:f^{-1}(Y_{1})\to Y_{1}$. Similarly to Theorem 1.1, we apply Theorem 4.5 putting $R+f^{*}S$ in the place of $R$ of Theorem 4.5 where $S\geq 0$ is an snc divisor supported on $B$ such that $R_{v1}+f^{*}S$ and $B_{R}+S\geq 0$. Let $N=O(S)$ be the associated $\mathbb{Q}$-line bundle.
We take the $L^{2}$ metric $\mu$ for $M\otimes f_{*}A$ (on $Y_{1,\text{free}}$) induced by a psh metric $\lambda$ for the $\mathbb{Q}$-line bundle $L+f^{*}N+A$ given by the effective divisor
$R+f^{*}S+\lceil-R_{h}\rceil=R_{h}+\lceil-R_{h}\rceil+R_{v1}+f^{*}S+R_{v2}.$
Now Theorem 4.5 (2) says that $\mu$ and $\varphi_{B_{R}+S}$ are valuatively equivalent where $\varphi_{B_{R}+S}$ has divisorial poles along the snc divisor $B_{R}+S$. Take the curvature currents: let $T$ be the curvature current of the psh metric (i.e. the singular hermitian line bundle) $(J+H+N,\mu)$ and $Q$ the curvature current of the psh metric $(H+N,\varphi_{B_{R}+S})$. From the Siu decomposition of $T$, we have $R_{T}$ with vanishing Lelong numbers, belonging to the first Chern class of the $\mathbb{Q}$-line bundle $J$. This gives rise to a psh metric $\psi$ with vanishing Lelong numbers for $J$. Subtracting $(J,\psi)$ from $(J+H+N,\mu)$, we obtain a psh metric $(H+N,\mu-\psi)$ which is one given by the effective divisor $B_{R}+S$.
Choose a singular hermitian metric $\varphi_{S}$ given by the divisor $S$. Subtracting again $(N,\varphi_{S})$, we get $(H,\varphi_{B_{R}}=\mu-\psi-\varphi_{S})$ which is a singular hermitian metric given by the original discriminant divisor $B_{R}$. Here we can and do arrange so that $\varphi_{B_{R}+S}=\varphi_{B_{R}}+\varphi_{S}$.
Step 3. Finally we use the fiberwise integration property of the $L^{2}$ metric $\mu$ on $Y_{1}$ given in Proposition 3.3 which continues to hold in the $\mathbb{Q}$-line bundle setting of Proposition 3.6.
Recall that in the notation of Proposition 3.3, $u$ and $\eta$ are local sections so that $\left|u(\eta)\right|^{2}$ and $\left|u\cdot\eta\right|^{2}$ are (local) smooth volume forms. The fiberwise integration of the singular volume form $\left|u(\eta)\right|^{2}e^{-\lambda}$ is equal to $\left|u\cdot\eta\right|^{2}e^{-\mu}$. Hence the fiberwise integration of $\tilde{\beta}:=\left|u(\eta)\right|^{2}e^{-(\lambda-f^{*}\varphi_{S})}$ is equal to $\left|u\cdot\eta\right|^{2}e^{-(\mu-\varphi_{S})}$, which we take as $\theta$. Note that $\tilde{\beta}$ has divisorial poles given by $e^{-(\lambda-f^{*}\varphi_{S})}$ which is along the divisor $R+\lceil-R_{h}\rceil$. It has equivalent singularities with the given $\beta$.
Since $\alpha$ has divisorial poles along the divisor $R$, we have $\alpha$ less singular than $\tilde{\beta}$ and thus $f_{*}\alpha$ less singular than $f_{*}\tilde{\beta}=\theta$. ∎
Remark 5.2.
In Theorem 1.2, we have ‘less singular’ instead of ‘equivalent singularities’ since a componenet of $R_{h}$ may be possibly not transversal with a special smooth fiber of $F$ lying on a point of $B$. (Note that $R_{h}$ is relative snc over $Y\setminus B$, cf. Definition 4.1.) Even though the coefficients of $R_{h}+\lceil-R_{h}\rceil$ are less than $1$, $(R_{h}+\lceil-R_{h}\rceil)|_{F}$ may not be an snc divisor and may possibly contribute to the difference of singularities between $f_{*}\alpha$ and $f_{*}\beta$. However, such contribution is rather mild in that it is still a locally integrable singularity on $Y$ due to the fact that both $R_{h}$ and $R_{h}+\lceil-R_{h}\rceil$ are klt divisors on $X$.
Remark 5.3.
Theorem 1.1 also yields a volume asymptotics from the fiberwise integration property of the $L^{2}$ metric. In the case when $\dim Y=1$, such a volume asymptotics can be compared with many previous results, especially recently by [BJ17], [EFM18, (2.1)] (see also related works by [Y10], [GTZ16, (2.1)], [GTZ19], [Be16, (3.8)] and others). Such computations for $\int_{X_{t}}\left|\sigma_{t}\right|^{2}$ can be roughly (and typically) summarized as (for a local analytic coordinate $t$ on $Y$ with $\dim Y=1$)
(18) $\left|t\right|^{-2\alpha}\left|\log\left|t\right|^{2}\right|^{\beta}$
where $\alpha\geq 0$ can be interpreted in terms of log-canonical thresholds.
6. Applications to $L^{2}$ extension theorems
Let $(X,\psi)$ be a pair where $X$ is a complex manifold and $\psi$ is a psh metric with neat analytic singularities for a $\mathbb{Q}$-line bundle $L$ on $X$. A log canonical center (or an lc center) of $(X,\psi)$ is an irreducible subvariety $Y\subset X$ that is the image of a prime divisor $E$ in a log resolution $X^{\prime}\to X$ of the pair, with its discrepancy $a(E,X,\psi)$ equal to $-1$ (i.e. the lowest possible value for an lc pair), cf. [Ko97, Def. 3.3], [K23, Def. 2.3]. We call such $E$ a log canonical place (or an lc place) of $Y$. When $Y$ is an lc center and there is no other lc center $Y_{1}$ such that $Y_{1}\supsetneq Y$, we call $Y$ a maximal lc center. The maximal lc centers are precisely the irreducible components of the non-klt locus of $(X,\psi)$. (We refer to [K23, Section 2] for more details on this setting.)
6.1. Ohsawa measure
Now let $Y$ be a maximal lc center of an lc pair $(X,\psi)$ as above. Let $\Psi$ be a quasi-psh function on $X$ with analytic singularities determined by the relation $e^{-\Psi}h=e^{-\psi}$ where $h$ is a smooth hermitian metric of $L$. In this setting, the Ohsawa measure $dV[\Psi]_{Y}$ is defined following the original definition of [O01], cf. [D15].
Definition 6.1.
[D15, Section 2] cf. [K23, Def. 3.1] Let $dV_{X}$ be a smooth volume form on $X$. The Ohsawa measure $dV[\Psi]_{Y}$ of $\Psi$ on $Y$ (with respect to $dV_{X}$) is a positive measure $d\mu$ on $Y_{\operatorname{reg}}$ satisfying the following condition: for every $g$, a real-valued compactly supported continuous function on $Y_{\operatorname{reg}}$ and for every $\tilde{g}$, a compactly supported extension of $g$ to $X$, we have the relation
[…]
See also [K23, Prop. 3.2] for the existence of this measure. Now assume that $Y$ has a unique lc place $E\to Y$ where $E\subset X^{\prime}$ is on a log resolution $\mu:X^{\prime}\to X$ of $(X,\psi)$. Then the morphism $E\to Y$ has connected fibers (cf. [K23, Cor. 2.12]). Writing the discrepancy relation as in [K23, (5) in Proof of Prop. 3.2], we have
(20) $K_{X^{\prime}}+E+F\equiv f^{*}(K_{X}+[\Psi]).$
By [K23, Prop. 3.2], the Ohsawa measure $dV[\Psi]_{Y}$ is equal to the direct image $(\mu|_{E})_{*}d\nu$ where $d\nu$ is a measure corresponding to a singular volume form with poles along the snc divisor $R:=F|_{E}$.
Following [Ka98], we may assume (by choosing $\mu:X^{\prime}\to X$) that the restriction $\mu|_{E}:E\to Y$ factors through a proper modification $\pi:Y^{\prime}\to Y$ and $f:(E,R)\to(Y^{\prime},B)$ satisfies Kawamata’s condition of Definition 4.1 for a suitable reduced snc divisor $B$ on $Y^{\prime}$. We have the following diagram.
[…]
Now applying Theorem 1.2 to $E\to Y^{\prime}$, we have the following description of the Ohsawa measure in terms of the discriminant divisor $B_{R}$ supported on $B$ as a consequence of Theorem 1.2.
Theorem 6.2.
Let $(X,\Psi)$ and $Y$ be as above, with a unique lc place as in (21). The Ohsawa measure $dV[\Psi]_{Y}$ is ‘less singular’ than the direct image $\pi_{*}v$ under $\pi:Y^{\prime}\to Y$ of a measure $v$ on $Y^{\prime}$ with poles along Kawamata’s discriminant divisor $B_{R}$ on $Y^{\prime}$ multiplied by a local psh weight with vanishing Lelong numbers. In other words, $v$ can be locally written as (up to a bounded positive factor)
(22) $v(w)=(\prod^{m}_{i=1}\left|w_{i}\right|^{-2a_{i}})e^{-\psi(w)}\left|dw_{1}\wedge\ldots\wedge dw_{m}\right|^{2}$
in local analytic coordinates $w=(w_{1},\ldots,w_{m})$ adapted to $B$ on an open subset $U\subset Y^{\prime}$ (with $\dim Y^{\prime}=\dim Y=m$) where $\psi$ is a psh function on $U$ with vanishing Lelong numbers and $\sum a_{i}\operatorname{div}(w_{i})=(B_{R})|_{U}$.
Proof.
As noted above, by [K23, Prop. 3.2], the Ohsawa measure $dV[\Psi]_{Y}$ is equal to the direct image $(\mu|_{E})_{*}d\nu$ where $d\nu$ is a measure corresponding to a singular volume form with poles along the snc divisor $R:=F|_{E}$.
Both $d\nu$ and $dV[\Psi]_{Y}$ have the property of putting no mass on closed analytic subsets, [K23, Prop. 3.2, Prop. 3.5]. Hence these measures are determined by their restrictions to nonempty Zariski open subsets. Therefore it suffices to check the property of ‘less singular’ on the restriction of $f:E\to Y^{\prime}$ to $f_{0}:f^{-1}(Y_{0})\to Y_{0}$ where $Y_{0}\subset Y^{\prime}$ is the subset of regular values of $f$.
Now it is a matter of applying Theorem 1.2 (for $f:X\to Y$ in the notation therein) to $f:E\to Y^{\prime}$ at hand. The conclusion follows from $f_{*}d\nu$ (‘the Ohsawa measure on $Y^{\prime}$’) being less singular than $v$, which results from $d\nu$ (resp. $\beta$) having the divisorial pole along $R=F|_{E}$ (resp. along $R+\lceil-R_{h}\rceil$) while $v$ at hand corresponds to $\theta=f_{*}\beta$ of Theorem 1.2. ∎
Remark 6.3.
Remark 5.2 applies also here. Note that even though $Y$ has a unique lc place, there may exist a component of $R_{h}$ which comes from a component of $F$ whose image under $\mu$ is equal to $Y$ but with discrepancy $>-1$.
6.2. Comparison of two $L^{2}$ extension theorems
Using Theorem 6.2, we establish a precise comparison between two $L^{2}$ extension theorems of [K10, Thm. 4.2] and [D15, Thm. 2.8] for log canonical pairs (for a smooth complex projective variety $X$). For this comparison, we first derive [K23, Thm. 3.9], a version of $L^{2}$ extension for a maximal log canonical center, from [D15, Thm. 2.8]. For the convenience of readers, we recall the statement.
Theorem 6.4.
[D15, Thm. 2.8 and Remark 2.9 (b)], cf. [K23, Thm. 3.9].
Let $(X,\omega)$ be a weakly pseudoconvex Kähler manifold. Let an lc pair $(X,\psi)$, a $\mathbb{Q}$-line bundle $L$, a psh metric $\psi$ for $L$ and a quasi-psh function $\Psi$ be as in Definition 6.1, so that $e^{-\psi}=he^{-\Psi}$. Assume that, for some $\delta>0$,
(23) $i\Theta(L,h)+\alpha\sqrt{-1}\partial\overline{\partial}\Psi=i\Theta(L,\psi)+(\alpha-1)\sqrt{-1}\partial\overline{\partial}\Psi\geq 0$
for all $\alpha\in[1,1+\delta]$. Let $B$ be a $\mathbb{Q}$-line bundle on $X$ such that $K_{X}+L+B$ is a $\mathbb{Z}$-line bundle. Let $b$ be a psh metric of $B$.
Let $Y$ be a maximal lc center of $(X,\psi)$ with a unique lc place. If a holomorphic section $s\in H^{0}(Y,(K_{X}+L+B)|_{Y})$ satisfies
[…]
then there exists a holomorphic section $\widetilde{s}\in H^{0}(X,K_{X}+L+B)$ such that we have $\tilde{s}|_{Y}=s$ and moreover
[…]
Here the Ohsawa measure $dV[\Psi]_{Y}$ is taken with respect to the smooth volume form $dV_{\omega}$ and $\gamma$ is as in [D15, (2.7)].
Let $\lVert s\rVert_{1}$ be the input norm of [K10, Thm. 4.2] which was given as the $L^{2}$ norm for an adjoint line bundle with respect to a Kawamata metric [K10, Def. 3.1]. Let $\lVert s\rVert_{2}$ be the input norm in [K23, Thm. 3.9] with respect to the Ohsawa measure $dV[\Psi]_{Y}$. From Theorem 6.2, we have the following comparison of two $L^{2}$ extension theorems from [K10] and [D15].
Theorem 6.5 (= Corollary 1.5).
Let $Y$ be a maximal lc center of $(X,\psi)$ with a unique lc place. Let $B$ be a $\mathbb{Q}$-line bundle on $X$ such that $K_{X}+L+B$ is a $\mathbb{Z}$-line bundle. Let $b$ be a psh metric of $B$. Let $s$ be a holomorphic section of $K_{Y}+M+B|_{Y}=(K_{X}+L+B)|_{Y}$ on $Y$ (where $K_{Y}+M$ is defined as in [K10, Def. 3.1]).
Let $p\in Y$ be a point. If the $L^{2}$ norm of $s$ with respect to a Kawamata metric $e^{-\kappa}$ is locally finite at $p$, i.e. there exists a neighborhood $V\subset Y$ of $p$ such that
$\lVert s\rVert_{1}:=\int_{V}\left|s\right|^{2}\cdot e^{-\kappa}\cdot b|_{Y}<\infty,$
then the $L^{2}$ norm of $s$ with respect to the Ohsawa measure is locally finite at $p$, i.e. there exists a neighborhood $U\subset Y$ of $p$ such that
[…]
Proof.
We will use the notation as in the setting of (21). In particular, we have the relation $K_{Y^{\prime}}+M^{\prime}=\pi^{*}(K_{Y}+M)$ where $\pi:Y^{\prime}\to Y$ is the modification as before.
From the definition of the Kawamata metric $e^{-\kappa}$ [K10, Def. 3.1], locally we can write $e^{-\kappa}=e^{-\varphi}e^{-\eta}$ as a metric for $M^{\prime}$ on $Y^{\prime}$ where $e^{-\varphi}$ is a singular hermitian metric given by the discriminant divisor and $e^{-\eta}$ is a choice of a smooth metric. On the other hand, by Theorem 6.2, the Ohsawa measure has density locally equal to the product $e^{-\varphi}e^{-\chi}$ where $\chi$ is a local psh weight with vanishing Lelong numbers. Let $\beta$ be a psh function such that $b=e^{-\beta}$ near $p$. Then by Lemma 6.6 (2), we have the equivalence:
$\int_{\pi^{-1}(V)}\left|\pi^{*}s\right|^{2}e^{-\varphi}e^{-\eta}e^{-\beta}<\infty\text{\;\; if and only if \;\;}\int_{\pi^{-1}(U)}\left|\pi^{*}s\right|^{2}e^{-\varphi}e^{-\chi}e^{-\beta}<\infty$
where $U$ and $V$ are as in the statement. This completes the proof. Note that $\varphi$ is locally the difference of two psh functions with analytic singularities, which is why we need (2) of Lemma 6.6. ∎
Lemma 6.6.
Let $u$ and $v$ be two psh functions on a complex manifold $X$. Suppose that $v$ has zero Lelong numbers at every point of $X$. Then the following hold:
(1) For the multiplier ideals, we have $\mathcal{J}(u+v)=\mathcal{J}(u)$.
(2) More generally, let $\psi$ be a psh function with analytic singularities. We have $e^{\psi-u}\in L^{2}_{\operatorname{loc}}$ if and only if $e^{\psi-u-v}\in L^{2}_{\operatorname{loc}}$.
Proof.
It is clear that (2) implies (1). (1) is from [K15, Prop. 2.3] and (2) from [B20, Cor. 10.15]. Both of them use the strong openness theorem of [GZ15]. ∎
7. Applications to semipositivity theorems
In this section, we will give an application of the main results to semipositivity theorems, Theorem 7.3. Before that, we first revisit the elliptic fibrations.
7.1. Elliptic fibrations
The classical elliptic fibrations studied by Kodaira [K63], [K64] provide the important initial case of the canonical bundle formula. Even in this case, our main result seems new, cf. Theorem 7.1.
Let $f:X\to Y$ be a relatively minimal elliptic fibration with $\dim Y=1$ (see e.g. [BPV, Thm. V.12.1]). When $f$ has multiple fibers $m_{1}F_{1},\ldots,m_{k}F_{k}$, we have the canonical bundle formula
(26) $K_{X}=f^{*}(K_{Y}+G)+\sum(m_{i}-1)F_{i}$
where $G$ is the line bundle equal to $f_{*}(K_{X/Y})$. We turn this into the following equality of $\mathbb{Q}$-line bundles:
(27) $K_{X}=f^{*}\left(K_{Y}+G+\sum^{k}_{i=1}\frac{m_{i}-1}{m_{i}}Q_{i}\right)$
where $Q_{i}\in Y$ is a point viewed as a divisor such that $f^{*}Q_{i}=m_{i}F_{i}$ as Cartier divisors. This is a situation where we can apply Theorem 3.7 viewing (27) as $K_{X}+L=f^{*}(K_{Y}+M)$ with $L$ trivial and $M:=G+\mathcal{O}(\sum^{k}_{i=1}\frac{m_{i}-1}{m_{i}}Q_{i})$. Note that in general, the equality $K_{X/Y}=f^{*}M$ holds as $\mathbb{Q}$-line bundles only : the direct image $f_{*}(K_{X/Y})$ is a $\mathbb{Z}$-line bundle which is not necessarily equal to $M$.
Furthermore, by [F86, (2.9)], [Ko07, (8.2.1)] (also see e.g. the introduction to [FM00]), we have equality of $\mathbb{Q}$-line bundles
(28) $G=\frac{1}{12}j^{*}\mathcal{O}_{\mathbf{P}^{1}}(1)+\mathcal{O}(\sum_{k\in K}\sigma_{k}P_{k})$
where $\sigma_{k}$ is the well-known coefficients (see e.g. [F86, (2.6)]) from the list of singular fibers [K63] and $j:Y\to\mathbf{P}^{1}$ is the map into the moduli.
The trivial metric $1$ is a psh metric for the trivial line bundle $L$, thus we get the corresponding $L^{2}$ metric $\mu$ for the $\mathbb{Q}$-line bundle $M$ in a canonical way.
Theorem 7.1.
In this case of an elliptic fibration $f:X\to Y$, the $L^{2}$ metric $\mu$ is the product of a singular hermitian metric given by the divisor $\sum_{k\in K}\sigma_{k}P_{k}+\sum^{k}_{i=1}\frac{m_{i}-1}{m_{i}}Q_{i}$ and a singular hermitian metric with vanishing Lelong numbers for $\frac{1}{12}j^{*}\mathcal{O}_{\mathbf{P}^{1}}(1)$.
Proof.
Let $g:X^{\prime}\to X$ be a proper birational morphism so that Kawamata’s condition (4.1) is satisfied for $X^{\prime}\to Y$ : we will apply Theorem 1.1. The $L^{2}$ metric for $X^{\prime}\to Y$ coincides with the $L^{2}$ metric $\mu$.
Define the divisor $R$ (in Theorem 1.1) by $K_{X^{\prime}}+R=g^{*}K_{X}$. For dimension reasons, $R_{h}=0$. The discriminant divisor $B_{R}$ is equal to $\sum_{k\in K}\sigma_{k}P_{k}+\sum^{k}_{i=1}\frac{m_{i}-1}{m_{i}}Q_{i}$. Thus as $\mathbb{Q}$-line bundles, $J$ (in Theorem 1.1) is equal to $\frac{1}{12}j^{*}\mathcal{O}_{\mathbf{P}^{1}}(1)$. This concludes the proof. ∎
Remark 7.2.
We note that already in this case, it is natural to adopt the generality of $\mathbb{Q}$-line bundles to be able to equally deal with multiple and non-multiple singular fibers in the canonical bundle formula, as in our results Theorem 1.1, Theorem 3.7. One can compare, for example, with [EFM18, Prop. 2.1] where the $L^{2}$ metric is for the line bundle $f_{*}(K_{X/Y})$ which seems to count non-multiple singular fibers only (as in (28)).
7.2. A question of Berndtsson on semipositivity theorems
We recall the following series of semipositivity theorems for an algebraic fiber space $f:X\to Y$ (i.e. a surjective morphism of smooth projective varieties with connected fibers) under some general conditions.
  1. (1)
    [Ka81, Thm. 5] : Nefness of the locally free sheaf $E:=f_{*}K_{X/Y}$. (Cf. [F78], [FFS].)
  2. (2)
    [Ka98, Thm. 2] : Log version of (1) for log Calabi-Yau fibrations $f$.
  3. (3)
    [Ka00, Thm. 1.1], [FF17] : Refinement of (1) replacing nefness by the existence of a singular hermitian metric with vanishing Lelong numbers for $\mathcal{O}_{\mathbf{P}(E)}(1)$.
We showed in Theorem 1.1 that the moduli part line bundle in the canonical bundle formula [Ka98, Thm. 2] for $f:X\to Y$ admits a psh metric with vanishing Lelong numbers when $R_{h}\geq 0$. In some cases, the moduli part can coincide with the direct image $f_{*}(K_{X/Y})$ of the relative canonical line bundle of $f$. Using this, we obtain as a corollary of Theorem 1.1, an alternative proof of the following Kawamata semipositivity theorem [Ka00, Thm. 1.1 (3)], [Ka81] when $f_{*}(K_{X/Y})$ is of rank $1$, which does not use some heavy machineries from the theory of variation of Hodge structure, cf. [CKS86], [S73].
Theorem 7.3.
[Ka00, Thm. 1.1 (3)], cf. [FF17, Cor. 1.2] Let $f:X\to Y$ be a surjective proper morphism with connected fibers between connected complex manifolds. Let $B$ be an snc divisor on $Y$ such that $f$ restricted over $Y\setminus B$ is a holomorphic submersion. Let $X_{0}:=f^{-1}(Y\setminus B)$. Let $n:=\dim X-\dim Y$.
  • Suppose that a general smooth fiber $F$ satisfies $K_{F}\sim 0$.
  • Suppose that $R^{n}f_{*}\mathbb{C}_{X_{0}}$ has unipotent monodromies around the components of $B$.
Then the line bundle $f_{*}(K_{X/Y})$ admits a singular hermitian metric with vanishing Lelong numbers.
The fact that $f_{*}(K_{X/Y})$ is a line bundle in this setting is due to [Ka81, §4] (see also e.g. [Ko07, (8.4.4)] for related exposition). This uses some well-known results from Hodge theory, for which we follow the exposition in [Ko07].
In [Ko07], they are used to define the moduli part line bundle $J(X/Y,R)$ and show the equality in (8.5.1) of $\mathbb{Q}$-line bundles $K_{X}+R=f^{*}(K_{Y}+J(X/Y,R)+B_{R})$, which is nontrivial this time unlike (11) since here $J(X/Y,R)$ has its own Hodge-theoretic definition unlike (11).
With this Hodge-theoretic characterization of the moduli part line bundle, one knows that it is equal to $f_{*}K_{X/Y}$ as in [Ko07, (8.4.4)] which is used in the following proof.
Proof.
First observe that, by [Ko07, Lem. 8.3.4], there exists a vertical divisor $R$ on $X$ such that $K_{X}+R$ is $\mathbb{Q}$-linearly equivalent to the pullback by $f$ of a $\mathbb{Q}$-line bundle on $Y$. Let $\mu:X^{\prime}\to X$ be a proper modification (given by composition of blow-ups) such that $f^{\prime}:X^{\prime}\to X\to Y$ satisfies Kawamata’s condition (4.1). Then one can write (for a divisor $R^{\prime}$ on $X^{\prime}$) that $K_{X^{\prime}}+R^{\prime}=\mu^{*}(K_{X}+R)=f^{\prime*}(K_{Y}+J(X^{\prime}/Y,R^{\prime})+B_{R^{\prime}})$ from [Ko07, Thm. 8.3.7].
As [Ko07, Thm. 8.3.7 (1)] states, the moduli part depends only on the generic fibers $F$ of $f$ and the pairs on them $(F,R_{h}|_{F})$. In the present case, $R$ and $R^{\prime}$ are vertical divisors, thus we have $R_{h}=0,R^{\prime}_{h}=0$. For a generic fiber $F^{\prime}$ of $X^{\prime}\to Y$, we have isomorphism $F^{\prime}\to F$ and in view of the pairs $(F,R_{h}|_{F})=(F,0),(F^{\prime},R^{\prime}_{h}|_{F}^{\prime})=(F^{\prime},0)$, we have
(29) $J(X^{\prime}/Y,R^{\prime})=J(X/Y,R)$
on $Y$ where both sides are defined in [Ko07, (8.4.6)].
Now considering the open set $X^{\prime}_{0}\subset X^{\prime}$ that is isomorphic to $X_{0}$ under the proper modification $X^{\prime}\to X$ (indeed $X^{\prime}\to X$ can be taken to satisfy this), we have $R^{n}f_{*}\mathbb{C}_{X_{0}}=R^{n}f^{\prime}_{*}\mathbb{C}_{X^{\prime}_{0}}$. Since the unipotent monodromies condition is satisfied for both of them, from [Ko07, (8.4.6)], we have $J(X^{\prime}/Y,R^{\prime})=f^{\prime}_{*}(K_{X^{\prime}/Y})$ and $J(X/Y,R)=f_{*}(K_{X/Y})$ respectively.
Combining with (29), we have $f_{*}(K_{X/Y})\cong J(X^{\prime}/Y,R^{\prime})$. Since $J(X^{\prime}/Y,R^{\prime})$ admits a psh metric with vanishing Lelong numbers $\lambda$ by Theorem 1.1, so does $f_{*}(K_{X/Y})$ by taking the image (call it $\lambda^{\prime}$) of $\lambda$ under the above isomorphism. Both $\lambda^{\prime}$ and $\lambda$ are the $L^{2}$ metrics. ∎